# AgentFly: Fine-tuning LLM Agents without Fine-tuning LLMs

### v1

@@ Abstract
In this paper, we introduce a novel learning paradigm for adaptive Large Language Model (LLM) agents that eliminates the need for fine-tuning the underlying LLMs. Existing approaches are often either rigid, relying on static, handcrafted reflection workflows, or computationally intensive, requiring gradient updates of LLM model parameters. In contrast, our method enables low-cost continual adaptation via memory-based online reinforcement learning. We formalise this as a Memory-augmented Markov Decision Process (M-MDP), equipped with a neural case-selection policy to guide action decisions. Past experiences are stored in an episodic memory, either differentiable or non-parametric. The policy is continually updated based on environmental feedback through a memory rewriting mechanism, whereas policy improvement is achieved through efficient memory reading (retrieval). We instantiate our agent model in the deep research setting, namely AgentFly, which attains top-1 on GAIA validation ($87.88\%$ Pass@33) and $79.40\%$ on the test set. It reaches $66.6\%$ F1 and $80.4\%$ PM on the DeepResearcher dataset, outperforming the state-of-the-art training-based method, while case-based memory adds $4.7\%$ to $9.6\%$ absolute points on out-of-distribution tasks. Our approach offers a scalable and efficient pathway for developing generalist LLM agents capable of continuous, real-time learning without gradient updates, advancing machine learning towards open-ended skill acquisition and deep research scenarios. The code is available at https://github.com/Agent-on-the-Fly/AgentFly.
(a) AgentFly vs. Baselines on GAIA validation and test sets.
(b) Ablation study of AgentFly across benchmarks.
@@
(d) AgentFly’s accuracy improvement on OOD datasets.
Figure 1: Overview of AgentFly evaluation across baselines, benchmarks, memory designs and generalisation.
@@ 1 Introduction
To this end, we introduce AgentFly, a non-parametric, learn-on-the-fly framework for CBR (Smyth and McClave, 2001, Hatalis et al., 2025), instantiated as a planner–executor architecture grounded in a memory-based Markov Decision Process (MDP). AgentFly comprises three principal components: (i) a planner, (ii) a tool-enabled executor, and (iii) a growing Case Bank that stores past trajectories as episodic memory. Instead of relying solely on the LLM’s parametric memory, which is fixed after training, online case-based reasoning in AgentFly is implemented by storing rich episodic traces.
Our experiments are conducted on $4$ benchmarks, where GAIA (Mialon et al., 2023) for long-horizon tool use, DeepResearcher (Zheng et al., 2025) for real-time web research, SimpleQA (Wei et al., 2024) for factual precision, and HLE (Phan et al., 2025) for long-tail academic reasoning. We use a planner–executor architecture with GPT-$4.1$ as the planner and o44-mini as the default executor (o33 for GAIA), instrumented with tools, namely AgentFly. We achieve top-1 on GAIA validation ($87.88\%$ Pass@33) and $79.40\%$ on the private test leaderboard, and it reaches $66.6\%$ F11 and $80.4\%$ PM on the DeepResearcher dataset, outperforming the state-of-the-art training-based system, while case-based memory adds $4.7$ to $9.6$ absolute points on out-of-distribution tasks and yields $95.0\%$ PM on SimpleQA. To our knowledge, we are the first to cast case-based continual learning for LLM agents, achieving the top-level performance on the GAIA benchmark, thereby providing a principled framework for continual adaptation of Deep Research agents.
@@ 2.2 Tool-augmented LLM
Language agents increasingly incorporate external tools to overcome context limitations and computational bottlenecks. prompt-based methods, including WebGPT (Nakano et al., 2021), embed tool calls directly in the generation trace. However, tackling long-horizon tasks are often required multi-hop tool calls. Therefore, recent works propose multi-agent pipelines, such as AutoGen (Wu et al., 2023), OWL (Camel-AI, 2025) and DeerFlow (ByteDance, 2025) to coordinate specialised agents via dialogue. To address long-horizon decision-making in dynamic, multi-turn interactions with external tool environments, Agentic Reinforcement Learning (Agentic RL) has emerged as a promising training paradigm. This approach shifts LLM training from static task-solving (e.g., math or code) to dynamic, agent–environment reasoning. Supervised Fine-tuning methods, including Toolformer (Schick et al., 2023), API-Bench (Li et al., 2023), and GRPO-based optimisation (Wang et al., 2025, Qian et al., 2025, Feng et al., 2025) teach models when and how to invoke APIs, but require costly retraining and often assume a fixed, small toolset (e.g., Code and Search). However, without explicit planning, deciding when and which tools to invoke remains a major bottleneck for long-horizon tasks. We model planning as a stateful MDP with explicit memory for past cases. By bringing case-based reasoning into planning, the executor is steered toward strategic tool calls and consistently strong performance.
@@ 4 Implementation: Deep Research Agent
We implement stateful prompt engineering via M-MDP methodology (§ 3) in Deep Research scenarios (Huang et al., 2025), where agents must solve complex, long-horizon tasks by iteratively interacting with their environment, invoking external tools, retrieving information from external sources, and processing heterogeneous data for dynamic reasoning. As illustrated in Figure 3, AgentFly alternates between two core stages: Case-Based Planning and Tool-Based Execution.
Figure 3: The architecture of AgentFly with parametric memory. AgentFly is instantiated as a planner–executor framework alternating between Case-Based Planning (Stage 1) and Tool-Based Execution (Stage 2). The planner is an LLM-based CBR agent enhanced by a Case Memory module that supports both Write, which records new cases and online refines the Q-function, and Read, which retrieves cases via the learned retrieval policy for adaptive case selection. The executor is an LLM-based MCP client that invokes external tools hosted on the MCP servers through the MCP protocol.
@@ 4.1 Framework
To address the challenges of long-horizon reasoning, AgentFly follows the plan-and-act paradigm (Erdogan et al., 2025), where the planner and executor operate in an alternating loop to iteratively advance task completion. For effective coordination, AgentFly integrates three memory modules:Case Memory (vectorised storage of prior cases for high-level planning), Subtask Memory (text-based storage of active subtasks and their results), and Tool Memory (text-based logs of tool interactions for each subtask).
@@
The execution stage is managed by an Executor, powered by a general-purpose LLM, which is responsible for executing each subtask as an autonomous episode (Sumers et al., 2023) using the MCP protocol. Unlike prior agents (Zheng et al., 2025, Weng et al., 2025), AgentFly’s executor supports rich reasoning and flexible tool composition. For each subtask, the executor consults tool memory, determines the appropriate tool invocation, and updates results.which operates as a Model Context Protocol (MCP) [^1] https://github.com/modelcontextprotocol client. The executor reads pending subtasks from the subtask memory, accesses relevant history from a Tool Memory (scoped per subtask), and determines whether to invoke an external tool or return a result. MCP serves as a standardized, model-agnostic interface, enabling flexible coordination with diverse external tools and data sources. By unifying access under a single protocol layer, AgentFly can seamlessly integrate dynamic reasoning and compositional tool use across multiple domains.
@@ 4.2 Case Memory Management
Non-Parametric Memory Retrieval. A cornerstone of AgentFly is its dynamically evolving Case Bank, which underpins its continual learning capability. At each planning step, this non-parametric memory module receives the task instruction and then retrieves relevant cases, comprising a mixture of successful and failed cases. This CBR method mirrors human analogical learning, where previously encountered outcomes shape decision-making (Aamodt and Plaza, 1994). Specifically, we retrieve the $\mathrm{K}$ nearest past cases from the case bank by computing the semantic similarity between the current state and past states. This design follows the mainstream CBR paradigm, which assumes that similar problems should have similar solutions (Wiratunga et al., 2024, Guo et al., 2025), thereby allowing the agent to prioritise cases whose historical contexts are most aligned with the current task. Formally, the Read operator of the non-parametric memory is defined as:
@@
Parametric Memory Retrieval. To empower the agent to selectively leverage high-utility cases to augment planning from past experiences, we design a differential memory mechanism in AgentFly via a parametric Q-function. When writing new cases to the case bank, the parametric method, in contrast to the non-parametric approach that merely appends the tuple as in Eq. (12), concurrently updates the Q-function online. Meanwhile, with CBR applied only for planning in AgentFly, the CBR planner can be simplified to a single-step setting instead of a multi-step M-MDP. This single-step setting collapses the TD target in Eq. (10) to the immediate reward, thereby simplifying the learning objective. Without bootstrapping, the updating reduces to a supervised learning paradigm, which avoids non-stationary targets. Therefore, we can train a parametric Q-function $Q(s,c;\theta)$ end-to-end, dispensing with the kernel-based estimation in Eq. (9). Accordingly, the single-step Q-learning loss can be formulated as:
@@ 4.3 Tool Usage
Besides the inherent requirement for long task execution sequences and multi-turn interactions, deep research tasks also place stringent demands on the atomic actions, which require the agent to be able to acquire external information and subsequently process, integrate, and analyse it. Thus, we design a suite of tools for AgentFly accessible via the MCP protocol, comprising modules for information retrieval such as search engines and web crawlers, as well as components for processing and analysing multimodal information, including video and image data, and files in various formats.
@@
Reasoning. The reasoning and analysis toolkit integrates code execution and mathematical computation to support robust, automated analysis within the AgentFly framework. The Code tool provides a sandboxed environment for writing, running, and managing code within a unified workspace. Users can create files, execute shell or Python commands, and inspect outputs – all within a persistent task directory. Python scripts are validated against a security whitelist to ensure safe execution, supporting commonly used libraries such as numpy, pandas, and torch. The workspace maintains state across steps, enabling iterative development. This agent is crucial for solving data analysis, automation, or dynamic code generation tasks. Complementing this, the Math tool handles fundamental arithmetic operations.
@@ 5.1 Datasets
To evaluate the general-purpose reasoning capabilities of AgentFly, we adopt the GAIA benchmark (Mialon et al., 2023), which comprises 450 non-trivial questions with unambiguous answers – 300 in the test set and 150 in the validation set. Each question requires varying levels of tool use and autonomous planning, and the dataset is stratified into three difficulty levels: Level 1: Requires approximately 5 steps using a single tool; Level 2: Requires 5–10 steps involving multiple tools; Level 3: Involves up to 50 steps with no restrictions on the number or type of tools. Each level includes a public validation split and a private test split with hidden ground-truth answers and metadata.
We further evaluate AgentFly on broader benchmarks compiled in DeepResearcher (Zheng et al., 2025), which draws from seven open-domain QA datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQ) (Joshi et al., 2017), HotpotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), MusiQue (Trivedi et al., 2022), Bamboogle (Press et al., 2022), and PopQA (Mallen et al., 2022). Each dataset contributes 512 examples, except Bamboogle, which provides 125 high-quality samples curated to minimise contamination and emphasise web-based synthesis.
@@ 5.3 Model Configurations
The Offline Executor setting refers to one static executor, removing the planner, case memory, and all external tools, so it reflects only raw parametric knowledge from LLMs. The Online Executor starts from that stripped-down baseline but reconnects the same executor to live search and other MCP tools, reflecting the value of real-time retrieval and tool execution. AgentFly (w/o CBR) keeps episodic memory disabled, allowing us to measure the extra gain delivered specifically by case-based reasoning.
@@ 5.4 Experimental Results
Table 2: Top results on the GAIA Leaderboard as of June 26, 2025, AgentFly achieves the Top-1 performance on the validation set and the test set in open-source agent frameworks.
Deep Researcher. We include this dataset to test real-time web research, evidence retrieval, cross-page synthesis, and multi-hop reasoning. As shown in Table 1, AgentFly augmented with MCP tools (e.g., search engine, browser) reaches an average $66.6\%$ F1 across the seven DeepResearcher benchmarks, nearly doubling the $37.7\%$ F1 of the CoT + RAG baseline. This demonstrates that real-time, online retrieval tools can rival or even exceed carefully curated static databases.
GAIA (Validation & Test). To assess robustness in long-horizon planning, tool orchestration, and execution, we employ the GAIA benchmark. AgentFly attains the top-$1$ ranking on the validation set and $4$th place on the test set, outperforming most existing agent frameworks (Table 2). Notably, it surpasses widely used open-source frameworks, including Manus (Liang et al., 2025), Aworld (Alibaba, 2025), and OWL (Camel-AI, 2025), on both validation and test sets.
For the GAIA validation evaluation, we initialise memory from scratch and iteratively store both successful and failed trajectories in the case bank over three iterations. Using GPT-4.1 as the planner and o3 as the executor, AgentFly achieves $87.88\%$ accuracy on the validation set. For the GAIA test set, performance is based solely on the case bank accumulated during validation, yielding an accuracy of $79.40\%$. Although AgentFly demonstrates strong overall performance, challenges remain for Level 3 tasks that require extended reasoning horizons and advanced tool coordination.
@@
Humanity’s Last Exam (HLE). To evaluate the frontier of human knowledge and the complex reasoning ability in long-tail, specialised domains, we include the HLE  [^9] https://scale.com/leaderboard/humanitys_last_exam. Using our planner executor architecture, with planner GPT-$4.1$ and executor o44-mini with tools, AgentFly attains $24.4\%$ PM, ranking second overall and within $0.92$ points of GPT-$5$ at $25.32\%$, while outperforming Gemini-$2.5$-Pro at $21.64\%$, o33-high at $20.32\%$, and o44-mini-high at $18.08\%$. These results demonstrate that continual learning through CBR effectively transforms episodic experiences into reusable knowledge, offering a complementary pathway to generalisation even in long-tail domains where conventional tool usage and retrieval methods struggle.
SimpleQA. To evaluate AgentFly’s reliability and robustness against hallucination in single-hop factual question answering, we employ the SimpleQA benchmark. As illustrated in Figure 4, AgentFly, implemented with a planner–executor framework (GPT-4.1 as the planner and o4-mini as the executor) augmented with tool use, achieves the highest accuracy among all baselines. Specifically, it reaches an accuracy of $95.0\%$, outperforming WebSailor ($93.5\%$), WebDancer ($90.5\%$), WebThinker ($77.5\%$), and DeepSeek-r1-React ($72.2\%$). These results demonstrate that AgentFly provides strong factual reliability and substantially mitigates hallucination on straightforward single-hop queries, establishing a new state-of-the-art over prior web-agent baselines.

### v2

# Memento: Fine-tuning LLM Agents without Fine-tuning LLMs

@@ Abstract
In this paper, we introduce a novel learning paradigm for Adaptive Large Language Model (LLM) agents that eliminates the need for fine-tuning the underlying LLMs. Existing approaches are often either rigid, relying on static, handcrafted reflection workflows, or computationally intensive, requiring gradient updates of LLM model parameters. In contrast, our method enables low-cost continual adaptation via memory-based online reinforcement learning. We formalise this as a Memory-augmented Markov Decision Process (M-MDP), equipped with a neural case-selection policy to guide action decisions. Past experiences are stored in an episodic memory, either differentiable or non-parametric. The policy is continually updated based on environmental feedback through a memory rewriting mechanism, whereas policy improvement is achieved through efficient memory reading (retrieval). We instantiate our agent model in the deep research setting, namely Memento, which attains top-1 on GAIA validation ($87.88\%$ Pass@33) and $79.40\%$ on the test set. It reaches $66.6\%$ F1 and $80.4\%$ PM on the DeepResearcher dataset, outperforming the state-of-the-art training-based method, while case-based memory adds $4.7\%$ to $9.6\%$ absolute points on out-of-distribution tasks. Our approach offers a scalable and efficient pathway for developing generalist LLM agents capable of continuous, real-time learning without gradient updates, advancing machine learning towards open-ended skill acquisition and deep research scenarios. The code is available at https://github.com/Agent-on-the-Fly/Memento.
(a) Memento vs. Baselines on GAIA validation and test sets.
(b) Ablation study of Memento across benchmarks.
@@
(d) Memento’s accuracy improvement on OOD datasets.
Figure 1: Overview of Memento evaluation across baselines, benchmarks, memory designs and generalisation.
@@ 1 Introduction
To this end, we introduce Memento, a non-parametric, learn-on-the-fly framework for CBR (Smyth and McClave, 2001, Hatalis et al., 2025), instantiated as a planner–executor architecture grounded in a memory-based Markov Decision Process (MDP). Memento comprises three principal components: (i) a planner, (ii) a tool-enabled executor, and (iii) a growing Case Bank that stores past trajectories as episodic memory. Instead of relying solely on the LLM’s parametric memory, which is fixed after training, online case-based reasoning in Memento is implemented by storing rich episodic traces.
Our experiments are conducted on $4$ benchmarks, where GAIA (Mialon et al., 2023) for long-horizon tool use, DeepResearcher (Zheng et al., 2025) for real-time web research, SimpleQA (Wei et al., 2024) for factual precision, and HLE (Phan et al., 2025) for long-tail academic reasoning. We use a planner–executor architecture with GPT-$4.1$ as the planner and o44-mini as the default executor (o33 for GAIA), instrumented with tools, namely Memento. We achieve top-1 on GAIA validation ($87.88\%$ Pass@33) and $79.40\%$ on the private test leaderboard, and it reaches $66.6\%$ F11 and $80.4\%$ PM on the DeepResearcher dataset, outperforming the state-of-the-art training-based system, while case-based memory adds $4.7$ to $9.6$ absolute points on out-of-distribution tasks and yields $95.0\%$ PM on SimpleQA. To our knowledge, we are the first to cast case-based continual learning for LLM agents, achieving the top-level performance on the GAIA benchmark, thereby providing a principled framework for continual adaptation of Deep Research agents.
@@ 2.2 Tool-augmented LLM
Language agents increasingly incorporate external tools to overcome context limitations and computational bottlenecks. Prompt-based methods, including WebGPT (Nakano et al., 2021), embed tool calls directly in the generation trace. However, tackling long-horizon tasks often requires multi-hop tool calls. Therefore, recent works propose multi-agent pipelines, such as AutoGen (Wu et al., 2023), OWL (Camel-AI, 2025) and DeerFlow (ByteDance, 2025) to coordinate specialised agents via dialogue. To address long-horizon decision-making in dynamic, multi-turn interactions with external tool environments, Agentic Reinforcement Learning (Agentic RL) has emerged as a promising training paradigm. This approach shifts LLM training from static task-solving (e.g., math or code) to dynamic, agent–environment reasoning. Supervised Fine-tuning methods, including Toolformer (Schick et al., 2023), API-Bench (Li et al., 2023), and GRPO-based optimisation (Wang et al., 2025, Qian et al., 2025, Feng et al., 2025) teach models when and how to invoke APIs, but require costly retraining and often assume a fixed, small toolset (e.g., Code and Search). However, without explicit planning, deciding when and which tools to invoke remains a major bottleneck for long-horizon tasks. We model planning as a stateful MDP with explicit memory for past cases. By bringing case-based reasoning into planning, the executor is steered toward strategic tool calls and consistently strong performance.
@@ 4 Implementation: Deep Research Agent
We implement stateful prompt engineering via M-MDP methodology (§ 3) in Deep Research scenarios (Huang et al., 2025), where agents must solve complex, long-horizon tasks by iteratively interacting with their environment, invoking external tools, retrieving information from external sources, and processing heterogeneous data for dynamic reasoning. As illustrated in Figure 3, Memento alternates between two core stages: Case-Based Planning and Tool-Based Execution.
Figure 3: The architecture of Memento with parametric memory. Memento is instantiated as a planner–executor framework alternating between Case-Based Planning (Stage 1) and Tool-Based Execution (Stage 2). The planner is an LLM-based CBR agent enhanced by a Case Memory module that supports both Write, which records new cases and online refines the Q-function, and Read, which retrieves cases via the learned retrieval policy for adaptive case selection. The executor is an LLM-based MCP client that invokes external tools hosted on the MCP servers through the MCP protocol.
@@ 4.1 Framework
To address the challenges of long-horizon reasoning, Memento follows the plan-and-act paradigm (Erdogan et al., 2025), where the planner and executor operate in an alternating loop to iteratively advance task completion. For effective coordination, Memento integrates three memory modules:Case Memory (vectorised storage of prior cases for high-level planning), Subtask Memory (text-based storage of active subtasks and their results), and Tool Memory (text-based logs of tool interactions for each subtask).
@@
The execution stage is managed by an Executor, powered by a general-purpose LLM, which is responsible for executing each subtask as an autonomous episode (Sumers et al., 2023) using the MCP protocol. Unlike prior agents (Zheng et al., 2025, Weng et al., 2025), Memento’s executor supports rich reasoning and flexible tool composition. For each subtask, the executor consults the tool memory, determines the appropriate tool invocation, and updates the results.which operates as a Model Context Protocol (MCP) [^1] https://github.com/modelcontextprotocol client. The executor reads pending subtasks from the subtask memory, accesses relevant history from a Tool Memory (scoped per subtask), and determines whether to invoke an external tool or return a result. MCP serves as a standardised, model-agnostic interface, enabling flexible coordination with diverse external tools and data sources. By unifying access under a single protocol layer, Memento can seamlessly integrate dynamic reasoning and compositional tool use across multiple domains.
@@ 4.2 Case Memory Management
Non-Parametric Memory Retrieval. A cornerstone of Memento is its dynamically evolving Case Bank, which underpins its continual learning capability. At each planning step, this non-parametric memory module receives the task instruction and then retrieves relevant cases, comprising a mixture of successful and failed cases. This CBR method mirrors human analogical learning, where previously encountered outcomes shape decision-making (Aamodt and Plaza, 1994). Specifically, we retrieve the $\mathrm{K}$ nearest past cases from the case bank by computing the semantic similarity between the current state and past states. This design follows the mainstream CBR paradigm, which assumes that similar problems should have similar solutions (Wiratunga et al., 2024, Guo et al., 2025), thereby allowing the agent to prioritise cases whose historical contexts are most aligned with the current task. Formally, the Read operator of the non-parametric memory is defined as:
@@
Parametric Memory Retrieval. To empower the agent to selectively leverage high-utility cases to augment planning from past experiences, we design a differential memory mechanism in Memento via a parametric Q-function. When writing new cases to the case bank, the parametric method, in contrast to the non-parametric approach that merely appends the tuple as in Eq. (12), concurrently updates the Q-function online. Meanwhile, with CBR applied only for planning in Memento, the CBR planner can be simplified to a single-step setting instead of a multi-step M-MDP. This single-step setting collapses the TD target in Eq. (10) to the immediate reward, thereby simplifying the learning objective. Without bootstrapping, the updating reduces to a supervised learning paradigm, which avoids non-stationary targets. Therefore, we can train a parametric Q-function $Q(s,c;\theta)$ end-to-end, dispensing with the kernel-based estimation in Eq. (9). Accordingly, the single-step Q-learning loss can be formulated as:
@@ 4.3 Tool Usage
Besides the inherent requirement for long task execution sequences and multi-turn interactions, deep research tasks also place stringent demands on the atomic actions, which require the agent to be able to acquire external information and subsequently process, integrate, and analyse it. Thus, we design a suite of tools for Memento accessible via the MCP protocol, comprising modules for information retrieval such as search engines and web crawlers, as well as components for processing and analysing multimodal information, including video and image data, and files in various formats.
@@
Reasoning. The reasoning and analysis toolkit integrates code execution and mathematical computation to support robust, automated analysis within the Memento framework. The Code tool provides a sandboxed environment for writing, running, and managing code within a unified workspace. Users can create files, execute shell or Python commands, and inspect outputs – all within a persistent task directory. Python scripts are validated against a security whitelist to ensure safe execution, supporting commonly used libraries such as numpy, pandas, and torch. The workspace maintains state across steps, enabling iterative development. This agent is crucial for solving data analysis, automation, or dynamic code generation tasks. Complementing this, the Math tool handles fundamental arithmetic operations.
@@ 5.1 Datasets
To evaluate the general-purpose reasoning capabilities of Memento, we adopt the GAIA benchmark (Mialon et al., 2023), which comprises 450 non-trivial questions with unambiguous answers – 300 in the test set and 150 in the validation set. Each question requires varying levels of tool use and autonomous planning, and the dataset is stratified into three difficulty levels: Level 1: Requires approximately 5 steps using a single tool; Level 2: Requires 5–10 steps involving multiple tools; Level 3: Involves up to 50 steps with no restrictions on the number or type of tools. Each level includes a public validation split and a private test split with hidden ground-truth answers and metadata.
We further evaluate Memento on broader benchmarks compiled in DeepResearcher (Zheng et al., 2025), which draws from seven open-domain QA datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQ) (Joshi et al., 2017), HotpotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), MusiQue (Trivedi et al., 2022), Bamboogle (Press et al., 2022), and PopQA (Mallen et al., 2022). Each dataset contributes 512 examples, except Bamboogle, which provides 125 high-quality samples curated to minimise contamination and emphasise web-based synthesis.
@@ 5.3 Model Configurations
The Offline Executor setting refers to one static executor, removing the planner, case memory, and all external tools, so it reflects only raw parametric knowledge from LLMs. The Online Executor starts from that stripped-down baseline but reconnects the same executor to live search and other MCP tools, reflecting the value of real-time retrieval and tool execution. Memento (w/o CBR) keeps episodic memory disabled, allowing us to measure the extra gain delivered specifically by case-based reasoning.
@@ 5.4 Experimental Results
Table 2: Top results on the GAIA Leaderboard as of June 26, 2025, Memento achieves the Top-1 performance on the validation set and the test set in open-source agent frameworks.
Deep Researcher. We include this dataset to test real-time web research, evidence retrieval, cross-page synthesis, and multi-hop reasoning. As shown in Table 1, Memento augmented with MCP tools (e.g., search engine, browser) reaches an average $66.6\%$ F1 across the seven DeepResearcher benchmarks, nearly doubling the $37.7\%$ F1 of the CoT + RAG baseline. This demonstrates that real-time, online retrieval tools can rival or even exceed carefully curated static databases.
GAIA (Validation & Test). To assess robustness in long-horizon planning, tool orchestration, and execution, we employ the GAIA benchmark. Memento attains the top-$1$ ranking on the validation set and $4$th place on the test set, outperforming most existing agent frameworks (Table 2). Notably, it surpasses widely used open-source frameworks, including Manus (Liang et al., 2025), Aworld (Alibaba, 2025), and OWL (Camel-AI, 2025), on both validation and test sets.
For the GAIA validation evaluation, we initialise memory from scratch and iteratively store both successful and failed trajectories in the case bank over three iterations. Using GPT-4.1 as the planner and o3 as the executor, Memento achieves $87.88\%$ accuracy on the validation set. For the GAIA test set, performance is based solely on the case bank accumulated during validation, yielding an accuracy of $79.40\%$. Although Memento demonstrates strong overall performance, challenges remain for Level 3 tasks that require extended reasoning horizons and advanced tool coordination.
@@
Humanity’s Last Exam (HLE). To evaluate the frontier of human knowledge and the complex reasoning ability in long-tail, specialised domains, we include the HLE  [^9] https://scale.com/leaderboard/humanitys_last_exam. Using our planner executor architecture, with planner GPT-$4.1$ and executor o44-mini with tools, Memento attains $24.4\%$ PM, ranking second overall and within $0.92$ points of GPT-$5$ at $25.32\%$, while outperforming Gemini-$2.5$-Pro at $21.64\%$, o33-high at $20.32\%$, and o44-mini-high at $18.08\%$. These results demonstrate that continual learning through CBR effectively transforms episodic experiences into reusable knowledge, offering a complementary pathway to generalisation even in long-tail domains where conventional tool usage and retrieval methods struggle.
SimpleQA. To evaluate Memento’s reliability and robustness against hallucination in single-hop factual question answering, we employ the SimpleQA benchmark. As illustrated in Figure 4, Memento, implemented with a planner–executor framework (GPT-4.1 as the planner and o4-mini as the executor) augmented with tool use, achieves the highest accuracy among all baselines. Specifically, it reaches an accuracy of $95.0\%$, outperforming WebSailor ($93.5\%$), WebDancer ($90.5\%$), WebThinker ($77.5\%$), and DeepSeek-r1-React ($72.2\%$). These results demonstrate that Memento provides strong factual reliability and substantially mitigates hallucination on straightforward single-hop queries, establishing a new state-of-the-art over prior web-agent baselines.